Dont tell me weve lost another bid! exclaimed Janice Hudson, president of Prime Products Inc. Im afraid so, replied Dob. Determine the amount of manufacturing overhead cost that would have been applied to the Hastings job. Manufacturing overhe4. Assume that it is customary in the industry to bid jobs at 140% of total manufacturing cost (direct materials, direct labob. Compute the underapplied or overapplied overhead for the year, assuming that departmental overhead rates are used. Cutting
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Prime Products
Answer 1 a
Calculation of Plant wide OH Rate Cutting Machining Assembly Total Note
Manufacturing overhead       572,400.00 848,000.00 106,000.00 1,526,400.00 A
Direct Labor       318,000.00 212,000.00 424,000.00       954,000.00 B
Predetermined OH Rate                   1.60 C=A/B
Answer 1 b
Allocated OH for Hastings Job Amount $
Direct Labor         36,800.00 D
Predetermined OH Rate                   1.60 See C
Allocated OH         58,880.00 E=C*D
Answer 2 a
Calculation of Departmental OH Rate Cutting Machining Assembly Note
Manufacturing overhead       572,400.00 848,000.00 106,000.00 F
Direct Labor       318,000.00 212,000.00 424,000.00 G
Departmental OH Rate                   1.80               4.00               0.25 H=F/G
Answer 2 b
Allocated OH for Hastings Job Cutting Machining Assembly Total
Direct Labor         12,500.00       4,100.00     20,200.00         36,800.00 I
Departmental OH Rate                   1.80               4.00               0.25 See H
Allocated OH         22,500.00     16,400.00       5,050.00         43,950.00 J=H*I
Answer 4 a
Bid price under plant wide OH rate Amount $
Direct Materials         31,700.00 K
Direct Labor         36,800.00 See I
Allocated OH         58,880.00 See E
Total manufacturing cost       127,380.00 L=K+E+I
Bid price       178,332.00 M= L*140%
Answer 4 b
Bid price under departmental OH rate Amount $
Direct Materials         31,700.00 See K
Direct Labor         36,800.00 See I
Allocated OH         43,950.00 See J
Total manufacturing cost       112,450.00 N=K+I+J
Bid price       157,430.00 O=N*140%
Answer 5 a Amount $
Actual Labor cost       931,200.00 P
Predetermined OH Rate                   1.60 See C
Overhead applied 1,489,920.00 Q=P*C
Actual overhead    1,600,300.00 R
Under applied       110,380.00 S=R-Q
Answer 5 b
Overhead applied Cutting Machining Assembly
Direct Labor       344,000.00 228,000.00 359,200.00 T
Departmental OH Rate                   1.80               4.00               0.25 See H
Overhead applied       619,200.00 912,000.00     89,800.00 U=T*H
Remarks Overhead applied Actual Overhead Difference Remarks
Cutting       619,200.00 602,000.00     17,200.00 Over applied
Machining       912,000.00 901,000.00     11,000.00 Over applied
Assembly         89,800.00     97,200.00     (7,400.00) Under applied